### v2

@@ 2 Dynamic of Non-ideal Analog Update
The following theorem provides estimation of the approximation error. It has been shown empirically that the response granularity can be made sufficiently small for updating (Rao et al., 2023; Sharma et al., 2024), implying $\Delta w_{\min}\ll\Delta w$. Therefore, we establish the error estimation of the approximation based on small response granularity condition.
Theorem 2.1 (Error from discrete pulse update).
The proof of Theorem 2.1 is deferred to Appendix D. In Theorem 2.1, $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-U_{q}(w,\Delta w)|$ is the error between the true update and the proposed dynamic, while $|{\tilde{U}}^{\operatorname{BL}}_{q}(w,s)-w|$ is the difference between original weight and the updated one. Theorem 2.1 shows that the proposed dynamic dominates the update, and the approximation error is negligible when $\Delta w$ is small, which holds as $\Delta w$ always includes a small learning rate in gradient-based training.
Takeaway. Theorem 2.1 enables us to discuss the impact of response functions directly without dealing with element-specific details like update granularity $\Delta w_{\min}$ and cycle variation $\sigma_{c}$. Response functions are the bridge between the resistive element level equation (pulse update (2)) and algorithm level equation (dynamic of Analog Update (4)).
Figure 3: Examples of response functions from Definition 2.2; $w^{\diamond}$ is the symmetric point.
Compact formulations of vector. The update (4) holds at each resistive element, while the model $W$ contains $N$ resistive elements with different response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$. We stack all the weights $w_{k}$ and expected increment $\Delta w_{k}$ together into vectors $W_{k},\Delta W_{k}\in{\mathbb{R}}^{D}$, where $k$ is the iteration index [^1] This paper adopts $w$ to represent the element of the weight $W_{k}$ without specifying its index. The notation makes the formulations more concise and indicates that all elements are updated in parallel.. Similarly, the response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are stacked into $Q_{+}(\cdot)$ and $Q_{-}(\cdot)$, respectively. Let the notation $U_{Q}(W_{k},\Delta W)$ on matrices $W_{k}$ and $\Delta W$ denote the element-wise operation on $W_{k}$ and $\Delta W$, i.e. $[U_{Q}(W_{k},\Delta W)]_{i}:=U_{[Q]_{i}}([W_{k}]_{i},[\Delta w]_{i}),\forall i\in[D]$ with $[D]:=\{1,2,\cdots,D\}$ denoting the index set. The element-wise update (4) can be expressed as $W_{k+1}=U_{Q}(W_{k},\Delta W_{k})$. Leveraging the symmetric decomposition (Gokmen & Haensch, 2020; Wu et al., 2024a), we decompose $Q_{-}(W)$ and $Q_{+}(W)$ into symmetric component $F(\cdot)$ and asymmetric component $G(\cdot)$
$F(W):=(Q_{-}(W)+Q_{+}(W))/2,$ (6)
$G(W):=(Q_{-}(W)-Q_{+}(W))/2,$ (7)
@@
$W_{k+1}=W_{k}+\Delta W_{k}\odot F(W_{k})-|\Delta W_{k}|\odot G(W_{k})$ (8)
where ||$\cdot$|| ⋅ | and $\odot$ represent the element-wise absolute value and multiplication, respectively.
Gradient-based training algorithms on AIMC hardware. In (8), the desired update $\Delta W_{k}$ varies based on different algorithms. Replacing $\Delta W_{k}$ with the stochastic gradient $\nabla f(W_{k};\xi_{k})$, we obtain the dynamic of Analog SGD as
$W_{k+1}=$ $\ W_{k}-\alpha\nabla f(W_{k};\xi_{k})\odot F(W_{k})$ (9)
$\ -\alpha|\nabla f(W_{k};\xi_{k})|\odot G(W_{k})$
where $\alpha>0$ is the learning rate and $\xi_{k}$ is the noise introduced at time $k$. The noise can come from both data sampling and analog noise, like thermal or shot noise.
Saturation. To avoid reaching arbitrarily high or low conductance states, resistive elements get saturated when they keep receiving the same pulses. Constrained by that, the conductance of the resistive element with $q_{+}(\cdot)$ and $q_{-}(\cdot)$ is bounded by a dynamic range $[\tau^{\min},\tau^{\max}]$ where $\tau^{\min}$ and $\tau^{\max}$ are the saturation points with zero responses, i.e. $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$. Near the saturation points, the asymmetric issue is significant, and thus, the update in one direction is suppressed, considerably impacting the convergence. On the contrary, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, the analog update behaves like a digital update. We refer to $w^{\diamond}$ as symmetric point. Symmetric points are typically located in the interior of the dynamic range and are far from saturation.
Stacking all $w^{\diamond}$ into a vector $W^{\diamond}\in{\mathbb{R}}$. Observe that the function $G(W)$ is large near the saturation points while almost zero around $W^{\diamond}$, implying it can reflect the degree of saturation. At the same time, $F(W)$ is the average of the response functions in two directions. As we will see in Sections 3.2 and 4, the ratios $\frac{G(W)}{\sqrt{F(W)}}$ plays a critical role in the convergence behaviors.
Response function class. Since the behavior of resistive elements is always governed by physical laws, we are interested in a class of response functions that reflect some physical properties. First, the conductance ends up increasing when receiving a positive pulse and vice versa, leading to positive response functions. On top of that, we also assume the response functions are differentiable (and hence continuous) for mathematical tractability. Taking them and conductance saturation into account, we are interested in the following response function class.
Definition 2.2 (Response function class).
$q_{+}(\cdot)$ and $q_{-}(\cdot)$ with dynamic range $[\tau^{\min},\tau^{\max}]$ satisfy
@@
Definition 2.2 covers a wide range of response functions. Figure 3 showcases three examples from the response functions class, including linear, non-linear monotonic, and even non-monotonic functions.
@@ 3 Implicit Penalty and Inexact Convergence of Analog SGD
As a critical impact from the response functions, an implicit penalty term is applied to the objective, leading to an asymptotic error in Analog SGD.
@@ 3.1 Implicit penalty
We first give an intuition through a situation where $W_{k}$ is already a critical point, i.e., ${\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]=0$. Recall that stochastic gradient descent on digital hardware (Digital SGD) is stable in expectation, i.e.
${\mathbb{E}}_{\xi_{k}}[W_{k+1}]=W_{k}-{\mathbb{E}}_{\xi_{k}}[\alpha\nabla f(W_{k};\xi_{k})]=W_{k}.$ (10)
However, this does not work for Analog SGD
$\ {\mathbb{E}}_{\xi_{k}}[W_{k+1}]=W_{k}-{\mathbb{E}}_{\xi_{k}}[\alpha\nabla f(W_{k};\xi_{k})\odot F(W_{k})$
$\ \hskip 60.00009pt-\alpha|\nabla f(W_{k};\xi_{k})|\odot G(W_{k})]$ (11)
@@
Consider a simplified version that the weight is a scalar ($D=1$) and the function $G(W)$ is strictly monotonically decreasing [^2] It happens when both $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are strictly monotonic. to help us gain intuition on the impact of the drfit in (3.1). Recall $G(W^{\diamond})=0$ at the symmetric point $W^{\diamond}$. $G(W)>0$ when $W>W^{\diamond}$ and $G(W)<0$ otherwise. Consequently, (3.1) indicates that ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]<W_{k}$ when $W_{k}>W^{\diamond}$ and ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]>W_{k}$ otherwise. Consequently, $W_{k}$ suffers from a drift tendency towards $W^{\diamond}$. In addition, it can be speculated that the penalty coefficient supposes to be proportional to the noise level since the drift is proportional to ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})|]$, which is the first moment of noise ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})-{\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]|]$ in essence.
The following theorem formalizes the implicit penalty effect. Before that, we define an accumulated asymmetric function $R_{c}(\cdot):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$, whose gradient is $R(W):=\frac{G(W)}{F(W)}$, i.e. $\nabla R_{c}(W)=R(W)=\frac{G(W)}{F(W)}$. Consequently, if $R(W)$ is strictly monotonic, $R_{c}(W)$ reaches its minimum at the symmetric point $W^{\diamond}$ where $R(W^{\diamond})=0$ so that it penalizes the weight away symmetric point.
Theorem 3.1 (Implicit penalty, short version).
@@
The full version of Theorem 3.1 and its proof are deferred to Appendix E. In Theorem 3.1, $R_{c}(W)$ plays the role of penalty to force the weight toward a symmetric point. As shown in Appendix E, $R_{c}(W)$ has a simple expression on linear response functions when $c_{\texttt{Lin}}=0$, leading (12) to
$\min_{W}~{}f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ (13)
which has the objective with $\ell_{2}$ regularization. In addition, the implicit penalty has a coefficient proportional to the noise level $\Sigma$ and inversely proportional to the dynamic range $\tau$. It implies that the implicit penalty becomes active only when gradients are noisy, and it is amplified when the noise is large.
Implicit penalty. When the gradient is noisy, an implicit penalty attracts Analog SGD toward symmetric points.
3.2 Inexact Convergence of Analog SGD
@@
Assumption 3.2 (Objective).
The objective $f(W)$ is $L$-smooth and is lower bounded by $f^{*}$
Assumption 3.3 (Unbiasness and bounded variance).
The sample $\{\xi_{k}:k\in[K]\}$ are independently and identically sampled from a distribution over times $k\in[K]$. Furthermore, the stochastic gradient is unbiased and has bounded variance, i.e., ${\mathbb{E}}_{\xi_{k}}[\nabla f(W_{k};\xi_{k})]=\nabla f(W_{k})$ and ${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}$.
Figure 4: The test accuracy curves for the model training on MNIST dataset under different $\tau$; (Left) FCN. (Right) CNN.
Assumption 3.23.3 are standard in non-convex optimization (Bottou et al., 2018; Wu et al., 2024a). Additionally, similar to the setting in Wu et al. (2024a), we also assume that the saturation degree is bounded, given by a lower bound of response functions.
Assumption 3.4 (Bounded saturation).
There exists a constant $H_{\min}>0$ such that $\min\{Q_{+}(W)\odot Q_{-}(W)\}\!>\!H_{\min}$.
Assumption 3.4 requires that $W_{k}$ remains far from saturation points, which is a mild assumption in actual training. This paper considers the average gradient square norm as the convergence metric, given by $E_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
Theorem 3.5 (Inexact convergence of Analog SGD).
Under Assumption 3.23.4, if the learning rate is set as $\alpha=O(1/\sqrt{K})$, it holds that
@@
The proof of Theorem 3.5 is deferred to Appendix F. Theorem 3.5 suggests that the convergence metric $E_{K}$ is upper bounded by two terms: the first term vanishes at a rate of $O(\sqrt{{\sigma^{2}}/{K}})$, which matches the Digital SGD’s convergence rate (Bottou et al., 2018) up to a constant; the second term contributes to the asymptotic error of Analog SGD, which does not vanish with the number of iterations $K$.
@@
4 Mitigating Implicit Penalty by Residual Learning: Tiki-Taka
@@
Residual learning. Our solution overlaps the algorithmic stationary point and physical symmetric point on the special point 00. Besides the main analog array, $W_{k}$, we maintain another array, $P_{k}$, whose stationary point is 00. A natural choice is the residual of the weight given by the solution of the following problem
$P^{*}(W)\in\operatornamewithlimits{arg\,min}_{P\in{\mathbb{R}}^{D}}~{}f(W+\gamma P)$ (15)
where $\gamma\geq 0$ is a mixing coefficient. Noticing that 00 is a solution of (15) if $W$ is already a solution of (1). Furthermore, the gradient of $f(W+\gamma P)$ with respect to $P$, given by $\nabla_{P}f(W+\gamma P)=\gamma\nabla f(W+\gamma P)$, is accessible with fair expense, enabling us to introduce a sequence $P_{k}$ to track the residual of $W_{k}$ by optimizing (15)
$P_{k+1}=$ $\ P_{k}-\alpha\nabla f(\bar{W}_{k};\xi_{k})\odot F(P_{k})$ (16)
$\ -\alpha|\nabla f(\bar{W}_{k};\xi_{k})|\odot G(P_{k}).$
Ideally, it holds that $P_{k}\approx P^{*}(W_{k})$ after (16), and hence it can be expected that $f(W_{k}+\gamma P_{k})\leq f(W_{k})$ so that one step of transfer leads to descent on $f(W_{k})$
$W_{k+1}=$ $\ W_{k}+\beta P_{k+1}\odot F(W_{k})-\beta|P_{k+1}|\odot G(W_{k})$ (17)
which solves the following problem by approximating the gradient $P^{*}(W)$ by $P_{k}$
$\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}~{}\|P^{*}(W)\|^{2}.$ (18)
The updates (16) and (17) are performed alternatively until convergence [^3] In principle, $F(\cdot)$ and $G(\cdot)$ vary from $W_{k}$ and $P_{k}$. We adopt the same notations for them just for convenience.. By solving a bilevel optimization problem (15) and (18) approximately, our solution recovers the update of Tiki-Taka Gokmen & Haensch (2020).
On the response functions side, it is naturally required to let zero be a symmetric point, i.e. $G(0)=0$, which can be implemented by zero-shifting technique (Kim et al., 2019) by subtracting a reference array.
Convergence properties of Tiki-Taka. We begin by analyzing the convergence of Tiki-Taka without considering the zero-shift first, enabling us to understand how the zero-shift response function benefit the convergence.
If the optimal point $W^{*}$ exists and is unique, the solution of (15) has a closed form $P^{*}(W):=\frac{W^{*}-W}{\gamma}$. At that time, the objective of (18) is equivalent to $\|W^{*}-W\|^{2}$. However, the solutions of (15) and (18) are non-unique in general, especially for non-convex objectives with multiple local minima. To ensure the existence and uniqueness of $W^{*}$, we assume the objective is strongly convex.
Assumption 4.1 ($\mu$-strong convexity).

### v3

@@ 2 Analog Training with Generic Response Functions
The update (3) holds at each resistive element, while the model $W$ contains $D$ resistive elements with different response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$. We stack all the weights $w_{k}$ and expected increment $\Delta w_{k}$ together into vectors $W_{k},\Delta W_{k}\in{\mathbb{R}}^{D}$, where $k$ is the iteration index Similarly, the response functions $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are stacked into $Q_{+}(\cdot)$ and $Q_{-}(\cdot)$, respectively. Let the notation $U_{Q}(W_{k},\Delta W)$ on matrices $W_{k}$ and $\Delta W$ denote the element-wise operation on $W_{k}$ and $\Delta W$, i.e. $U_{Q}(W_{k},\Delta W)]_{i}:=U_{[Q]_{i}}([W_{k}]_{i},[\Delta w]_{i}),\forall i\in[D]$ with $D]:=\{1,2,\cdots,D\}$ denoting the index set. The element-wise update (3) can be expressed as $W_{k+1}=U_{Q}(W_{k},\Delta W_{k})$. Leveraging the symmetric decomposition [22, 21], we decompose $Q_{-}(W)$ and $Q_{+}(W)$ into symmetric component $F(\cdot)$ and asymmetric component $G(\cdot)$
$F(W):=(Q_{-}(W)+Q_{+}(W))/2,\quad\text{and}\quad G(W):=(Q_{-}(W)-Q_{+}(W))/2,$ (4)
@@
$\texttt{Analog~Update}\qquad W_{k+1}=W_{k}+\Delta W_{k}\odot F(W_{k})-|\Delta W_{k}|\odot G(W_{k}).$ (5)
Gradient-based training algorithms on AIMC hardware. In (5), the desired update $\Delta W_{k}$ varies based on different algorithms. Replacing $\Delta W_{k}$ with the stochastic gradient $\nabla f(W_{k};\xi_{k})$, we obtain the dynamics of Analog SGD shown in (2). This update is reduced to the mathematical form for linear response functions in [21] as a special case; see Appendix B for details. Response function class. Before proceeding to the study of response functions, we first define the response function class that interests us. Since the behavior of resistive elements is always governed by physical laws, the function class should reflect certain crucial physical properties.
We first introduce a commonly observed saturation property across a wide range of resistive elements. Resistive elements get saturated when they keep receiving the same pulses to avoid reaching arbitrarily high or low conductance states. Constrained by that, the conductance of the resistive element with $q_{+}(\cdot)$ and $q_{-}(\cdot)$ is bounded inside a dynamic range $\tau^{\min},\tau^{\max}]$ where $\tau^{\min}$ and $\tau^{\max}$ are the saturation points with zero responses, i.e. $q_{+}(\tau^{\max})=q_{-}(\tau^{\min})=0$. Near the saturation points, the asymmetric issue is significant, and thus, the update in one direction is suppressed, considerably impacting the convergence. On the contrary, if a point $w^{\diamond}$ satisfies $q_{-}(w^{\diamond})=q_{+}(w^{\diamond})$, the analog update behaves like a digital update. We refer to $w^{\diamond}$ as symmetric point. Symmetric points are typically located in the interior of the dynamic range and are far from saturation.
Stacking all $w^{\diamond}$ into a vector $W^{\diamond}\in{\mathbb{R}}^{D}$. Observe that the function $G(W)$ is large near the saturation points while almost zero around $W^{\diamond}$, implying it can reflect the saturation degree. At the same time, $F(W)$ is the average of the response functions in two directions. As we will see in Sections 3.2 and 4, the ratio $\frac{G(W)}{\sqrt{F(W)}}$ plays a critical role in the convergence behaviors.
Figure 3: Examples of response functions from Definition˜1; $w^{\diamond}$ is the symmetric point.
Besides saturation, the function class should also enjoy other properties. First, the conductance ends up increasing when receiving a positive pulse and vice versa, leading to positive response functions. On top of that, we also assume the response functions are differentiable (and hence continuous) for mathematical tractability. Taking all into account, we define the following response function class.
Definition 1 (Response function class).
$q_{+}(\cdot)$ and $q_{-}(\cdot)$ with dynamic range $\tau^{\min},\tau^{\max}]$ satisfy
@@
Definition˜1 covers a wide range of response functions, including but not limited to PCM, ReRAM, ECRAM, and others mention in Appendix˜A. Figure 3 showcases three examples from the response functions class, including linear, non-linear but monotonic, and even non-monotonic functions.
@@ 3 Implicit Penalty and Inexact Convergence of Analog SGD
This section introduces a critical impact of the response functions, implicit penalized objective. Affected by it Analog SGD can only converge inexactly with a non-diminishing asymptotic error.
@@ 3.1 Implicit penalty
We first give an intuition through a situation where $W_{k}$ is already a critical point, i.e., ${\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]=0$. Recall that stochastic gradient descent on digital hardware (Digital SGD) is stable in expectation, i.e. ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]=W_{k}-{\mathbb{E}}_{\xi_{k}}[\alpha\nabla f(W_{k};\xi_{k})]=W_{k}$. However, this does not work for Analog SGD
$\ {\mathbb{E}}_{\xi_{k}}[W_{k+1}]=W_{k}-{\mathbb{E}}_{\xi_{k}}[\alpha\nabla f(W_{k};\xi_{k})\odot F(W_{k})-\alpha|\nabla f(W_{k};\xi_{k})|\odot G(W_{k})]$ (6)
@@
Consider a simplified version that the weight is a scalar ($D=1$) and the function $G(W)$ is strictly monotonically decreasing [^1] It happens when both $q_{+}(\cdot)$ and $q_{-}(\cdot)$ are strictly monotonic. to help us gain intuition on the impact of the drift in (6). Recall $G(W^{\diamond})=0$ at the symmetric point $W^{\diamond}$. $G(W)>0$ when $W>W^{\diamond}$ and $G(W)<0$ otherwise. Consequently, (6) indicates that ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]<W_{k}$ when $W_{k}>W^{\diamond}$ and ${\mathbb{E}}_{\xi_{k}}[W_{k+1}]>W_{k}$ otherwise. It implies that $W_{k}$ suffers from a drift tendency towards $W^{\diamond}$. In addition, the penalty coefficient proportional to the noise level since the drift is proportional to ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})|]$, which is the first moment of noise ${\mathbb{E}}_{\xi_{k}}[|\nabla f(W_{k};\xi_{k})-{\mathbb{E}}_{\xi}[\nabla f(W_{k};\xi)]|]$ in essence.
The following theorem formalizes the implicit penalty effect. Before that, we define an accumulated asymmetric function $R_{c}(\cdot):{\mathbb{R}}^{D}\to{\mathbb{R}}^{D}$, whose derivative is $R(W):=\frac{G(W)}{F(W)}$, i.e. $\frac{{\mathrm{d}[R_{c}(W)]_{d}}}{{\mathrm{d}[W]_{d}}}=[R(W)]_{d}=\frac{[G(W)]_{d}}{[F(W)]_{d}}$. If $R(W)$ is strictly monotonic, $R_{c}(W)$ reaches its minimum at the symmetric point $W^{\diamond}$ where $R(W^{\diamond})=0$, so that it penalizes the weight away from the symmetric point.
Theorem 1 (Implicit penalty, short version).
@@
The full version of Theorem 1 and its proof are deferred to Appendix G. In Theorem 1, $R_{c}(W)$ plays the role of a penalty to force the weight toward a symmetric point. As shown in Appendix G, $R_{c}(W)$ has a simple expression on linear response functions when $c_{\texttt{Lin}}=0$, leading (7) to $\min_{W}~f_{\Sigma}(W):=f(W)+\frac{\Sigma}{2\tau}\|W\|^{2}$ which is an $\ell_{2}$ regularized objective. In addition, the implicit penalty has a coefficient proportional to the noise level $\Sigma$ and inversely proportional to the dynamic range $\tau$. It implies that the implicit penalty becomes active only when gradients are noisy, and the noise amplifies the effect.
With noisy gradients, an implicit penalty attracts Analog SGD towards symmetric points.
3.2 Inexact Convergence of Analog SGD under generic devices
@@
Assumption 1 (Objective).
The objective $f(W)$ is $L$-smooth and is lower bounded by $f^{*}$.
Assumption 2 (Unbiasness and bounded variance).
The stochastic gradient is unbiased and has bounded variance $\sigma^{2}$.
Assumption 12 are standard in non-convex optimization [38, 21]. Additionally, similar to the setting in [21], we also assume that the saturation degree is bounded, i.e., all response functions are positive.
Assumption 3 (Bounded saturation).
There exists a constant $R_{\min}>0$ such that $\min\{Q_{+}(W)\odot Q_{-}(W)\}\!>\!R_{\min}$.
Assumption 3 requires that $W_{k}$ remains far from saturation points, which is a mild assumption in actual training. This paper considers the average gradient square norm as the convergence metric, given by $E_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
Theorem 2 (Inexact convergence of Analog SGD).
Under Assumption 13, if the learning rate is set as $\alpha=O(1/\sqrt{K})$, it holds that
@@
The proof of Theorem 2 is deferred to Appendix H. Theorem 2 suggests that the convergence metric $E_{K}$ is upper bounded by two terms: the first term vanishes at a rate of $O(\sqrt{{\sigma^{2}}/{K}})$, which matches the Digital SGD’s convergence rate [38] up to a constant; the second term contributes to the asymptotic error of Analog SGD, which does not vanish with the number of iterations $K$.
@@
4 Mitigating Implicit Penalty by Residual Learning
@@
Residual learning. Our solution overlaps the algorithmic stationary point and physical symmetric point on the special point 0. Besides the main analog array, $W_{k}$, we maintain another array, $P_{k}$, whose stationary point should be 0. A natural choice is the residual of the weight, $P^{*}(W)$, defined by the $P$ that minimizes the objective $f(W+\gamma P)$ with a non-zero $\gamma$. Additionally, the goal of the main array is to minimize the residual so that the model $W_{k}$ is approaching optimal. This process can be formulated as the following bilevel problem whose optimal points can be proved to be that of $f(W)$
$\textsf{Residual Learning}\quad\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}~\|P^{*}(W)\|^{2},\quad\operatornamewithlimits{s.t.}~P^{*}(W)\in\operatornamewithlimits{arg\,min}_{P\in{\mathbb{R}}^{D}}~f(W+\gamma P).$ (9)
Now we propose a gradient-based method to solve (9). The gradient of $f(W+\gamma P)$ with respect to $P$, given by $\nabla_{P}f(W+\gamma P)=\gamma\nabla f(W+\gamma P)$, is accessible with fair expense, enabling us to introduce a sequence $P_{k}$ to track the residual of $W_{k}$ by optimizing $f(W_{k}+\gamma P)$
$P_{k+1}=$ $\ P_{k}-\alpha\nabla f(\bar{W}_{k};\xi_{k})\odot F(P_{k})-\alpha|\nabla f(\bar{W}_{k};\xi_{k})|\odot G(P_{k}).$ (10)
where $\bar{W}_{k}:=W_{k}+\gamma P_{k}$ is the mixed weight. We then derive the hyper-gradient of the upper-level objective. Notice $\nabla\|P^{*}(W)\|^{2}=2\nabla P^{*}(W)P^{*}(W)$. Assuming $W^{*}$ is the unique minimum of $f(\cdot)$, we know $P^{*}(W)$ satisfies $P^{*}(W)+W=W^{*}$. Taking gradient with respective to $W$ on both sides, we have $\nabla P^{*}(W)=-I$ and hence $\nabla\|P^{*}(W)\|^{2}=-2P^{*}(W)$. Approximating $P^{*}(W)$ by $P_{k}$, we reach the update of the main array.
$W_{k+1}=$ $\ W_{k}+\beta P_{k+1}\odot F(W_{k})-\beta|P_{k+1}|\odot G(W_{k}).$ (11)
Featuring moving the residual $P_{k}$ to $W_{k}$, (11) is referred to as transfer process. The updates (10) and (11) are performed alternatively until convergence. It is noticeable that Residual Learning has the same recursion format as Tiki-Taka. However, as we will discuss in Section˜5, the proposed method can be naturally extended, and simulations in Section˜6 show that the extension of Residual Learning outperforms the extension of Tiki-Taka.
On the response functions side, it is naturally required to let zero be a symmetric point, i.e., $G(0)=0$, which can be implemented by the zero-shifting technique [39] by subtracting a reference array.
Convergence properties of Residual Learning. We begin by analyzing the convergence of Residual Learningwithout considering the zero-shift first, which enables us to understand how zero-shifted response functions affect convergence.
If the optimal point $W^{*}$ exists and is unique, the solution of the lower-level objective has a closed form $P^{*}(W):=\frac{W^{*}-W}{\gamma}$. At that time, the upper-level objective equals $\|W^{*}-W\|^{2}$. However, the solutions of $f(\cdot)$ are non-unique in general, especially for non-convex objectives with multiple local minima. To ensure the existence and uniqueness of $W^{*}$, we assume the objective is strongly convex.
Assumption 4 ($\mu$-strong convexity).